# Generalized Method of Moments

### v1

@@ 2 Setup and the identified set
    The distribution $\pi$ of $(Z_{1},Z_{2})$ belongs to the set $\mathcal{P}(\mathcal{Z}_{1}\times\mathcal{Z}_{2})$ of absolutely continuous distributions with respect to a known probability measure $\mu$ on $\mathcal{Z}_{1}\times\mathcal{Z}_{2}\subset\mathbb{R}^{2d}$, where $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ are known compact sets.
@@
Assumptions 1(i), 1(iv) and 1(iii) are standard and impose compactness of the parameter space and the continuity of moments. Assumption 1(ii) is the requirement that the researcher knows which variables in the data are continuous and which are discrete. This knowledge determines the dominating measure $\mu$. Moreover, the researcher is assumed to be able to come up with a compact set $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ containing the support of the data. We impose no restrictions on the data-generating process beyond these basic requirements.
@@ 6 Monte Carlo simulation
Figure 2: Estimates of the identified set and confidence regions averaged over 10001000 simulations.
The simulation results are illustrated in Figure 2. The true identified set (red contour) is approximated using a sample of $10,000$ observations. Our estimates of the identified set track the true identified set closely. The confidence regions are based on $1000$ bootstrap samples and are indicated by three black contour lines, one for each nominal coverage $1-\alpha\in\{0.10,0.90,0.95\}$. They are found to be quite narrow even at high nominal confidence levels.
@@ Appendix B Proof of Proposition 2
Denote $c(z_{1},z_{2})=u^{\prime}\phi(z_{1},z_{2},\theta)$. Since $\pi^{0}\ll\mu$, each $\pi^{0}\in\Pi^{0}$ has a density $h=d\pi^{0}/d\mu$. Let
$\mu(dz_{1},dz_{2})=\mu_{1}(dz_{1})\,\mu(dz_{2}\mid z_{1})$
be the regular conditional decomposition of $\mu$. Then
$\pi^{0}(dz_{1},dz_{2})=h(z_{1},z_{2})\,\mu_{1}(dz_{1})\mu(dz_{2}\mid z_{1})$
and
$\int h(z_{1},z_{2})\,\mu(dz_{2}\mid z_{1})=\pi_{1}^{0}(z_{1})\qquad\text{for $\mu_{1}$-a.e.\ }z_{1}.$
For each $z_{1}$ with $\pi_{1}^{0}(z_{1})>0$, define
$g(z_{2}\mid z_{1}):=\frac{h(z_{1},z_{2})}{\pi_{1}^{0}(z_{1})}.$
Then $g(\cdot\mid z_{1})\geq 0$ and $\int g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})=1$, so $g(\cdot\mid z_{1})$ is a density with respect to $\mu(\cdot\mid z_{1})$. Therefore
$\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int\pi_{1}^{0}(z_{1})\left(\int c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\right)\mu_{1}(dz_{1}).$
Since any measurable family of densities $g(\cdot\mid z_{1})$ induces some $\pi^{0}\in\Pi^{0}$, maximizing over $\Pi^{0}$ reduces to choosing, for each $z_{1}$, a density $g(\cdot\mid z_{1})$ to maximize
$\int_{\mathcal{Z}_{2}}c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1}).$
Fix $z_{1}$ and write $h_{z_{1}}(z_{2}):=c(z_{1},z_{2})$. Since $h_{z_{1}}$ is continuous on the compact set $\mathcal{Z}_{2}$, it attains its maximum at some $z_{2}^{\ast}(z_{1})$. For any density $g(\cdot\mid z_{1})$,
$\int h_{z_{1}}(z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\leq h_{z_{1}}(z_{2}^{\ast}(z_{1})).$
Moreover, for any $\varepsilon>0$ the continuity of $h_{z_{1}}$ implies the existence of an open neighborhood $U_{\varepsilon}$ of $z_{2}^{\ast}(z_{1})$ such that $h_{z_{1}}(z_{2})\geq h_{z_{2}^{\ast}(z_{1})}-\varepsilon$ for all $z_{2}\in U_{\varepsilon}$, and $\mu(U_{\varepsilon}\mid z_{1})>0$. Normalizing the indicator of $U_{\varepsilon}$ yields a density $g_{\varepsilon}(\cdot\mid z_{1})$ for which
$\int h_{z_{1}}(z_{2})\,g_{\varepsilon}(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\geq h_{z_{1}}(z_{2}^{\ast}(z_{1}))-\varepsilon.$
Letting $\varepsilon\downarrow 0$ shows that the supremum over densities equals the maximum value $h_{z_{1}}(z_{2}^{\ast}(z_{1}))$:
$\sup_{g(\cdot\mid z_{1})}\int c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})=\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2}).$
Substituting this expression above yields
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int\pi_{1}^{0}(z_{1})\,\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\mu_{1}(dz_{1})=\operatorname{\mathbb{E}}\left[\max_{z_{2}\in\mathcal{Z}_{2}}c(Z_{1},z_{2})\,\middle|\,S=0\right],$
completing the proof.

### v2

@@ 2 Setup and the identified set
    The distribution $\pi$ of $(Z_{1},Z_{2})$ belongs to the set $\mathcal{P}(\mathcal{Z}_{1}\times\mathcal{Z}_{2})$ of probability distributions on $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, where $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ are known compact sets.
@@
Assumptions 1(i), 1(iv) and 1(iii) are standard and impose compactness of the parameter space and the continuity of moments. Assumption 1(ii) posits that the researcher knows the compact set $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ that contains (but does not have to be equal to) the true support of the data. We impose no restrictions on the data-generating process beyond these basic requirements.
@@ 4 Inference
Remark 4.2.
Using this uniform convergence result, it should be possible to develop a specification test based on the statistic $\max_{\theta\in\Theta}\hat{T}(\theta)$, the population analog of which is zero if and only if the identified set $\Theta_{I}$ is nonempty. The critical values can be obtained using the bootstrap of fang2019inference, in analogy to our treatment of the test statistic $\hat{T}(\theta_{0})$ for a fixed $\theta_{0}$ below. Similar specification tests have been developed for moment inequality models, see, e.g., bugni2015specification and references therein.
@@ 6 Monte Carlo simulation
Figure 2: Estimates of the identified set and confidence regions with nominal sizes $1-\alpha\in\{0.10,0.90,0.95\}$ averaged over 10001000 simulations.
The simulation results are illustrated in Figure 2. The true identified set (red contour line) is approximated using a sample of $10,000$ observations. Our estimates of the identified set track the true identified set closely. The confidence regions (three black contour lines) collect the hypothesized parameter values that are not rejected by the bootstrap test described in Section 4 with the size $\alpha\in\{0.05,0.10,0.90\}$ and the number of bootstrap samples $B=1000$. They are found to be quite narrow even at high nominal confidence levels $1-\alpha$.
@@ Appendix B Proof of Proposition 2
Denote $c(z_{1},z_{2})=u^{\prime}\phi(z_{1},z_{2},\theta)$ and let $\Pi^{0}$ be the set of probability measures $\pi^{0}$ on $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ with the first marginal $\pi_{1}^{0}(\cdot)=\mathbb{P}(Z_{1}\in\cdot\mid S=0)$.
Fix any $\pi^{0}\in\Pi^{0}$. There exists a Markov kernel $\pi^{0}(dz_{2}\mid z_{1})$ such that
$\pi^{0}(dz_{1},dz_{2})=\pi_{1}^{0}(dz_{1})\,\pi^{0}(dz_{2}\mid z_{1}),$
see, e.g., Theorems 10.2.1 and 10.2.2 in dudley2018real. Hence
$\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int_{\mathcal{Z}_{1}}\left(\int_{\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi^{0}(dz_{2}\mid z_{1})\right)\pi_{1}^{0}(dz_{1}).$
For each $z_{1}$,
$\int_{\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi^{0}(dz_{2}\mid z_{1})\leq\sup_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})=\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2}),$
where the equality follows from continuity of $c(z_{1},\cdot)$ and compactness of $\mathcal{Z}_{2}$. Integrating and taking the maximum over $\pi^{0}\in\Pi^{0}$ yields
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]\leq\int\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi_{1}^{0}(dz_{1}).$ (7)
We now establish a matching lower bound. Let $\{q(k)\}_{k\geq 1}$ be a countable dense subset of $\mathcal{Z}_{2}$. For each $N\geq 1$, define
$m_{N}(z_{1}):=\max_{1\leq k\leq N}c(z_{1},q(k))$
and let $k_{N}(z_{1})$ be an arbitrary element of the set
$\operatorname*{argmax}_{1\leq k\leq N}c(z_{1},q(k))=\{k\in\{1,\dots,N\}:\,\,c(z_{1},q(k))=m_{N}(z_{1})\}.$
Denote $q_{N}(z_{1})=q(k_{N}(z_{1}))$ so that $m_{N}(z_{1})=c(z_{1},q_{N}(z_{1}))$. Let $\delta_{q_{N}(z_{1})}(dz_{2})$ be the point mass at $q_{N}(z_{1})$. Define the joint distribution
$\pi_{N}(dz_{1},dz_{2}):=\pi_{1}^{0}(dz_{1})\,\delta_{q_{N}(z_{1})}(dz_{2}).$
Then $\pi_{N}\in\Pi^{0}$ and
$\operatorname{\mathbb{E}}_{\pi_{N}}[c(Z_{1},Z_{2})]=\int m_{N}(z_{1})\,\pi_{1}^{0}(dz_{1}).$
Since $\{q(k)\}_{k\geq 1}$ is dense in the compact set $\mathcal{Z}_{2}$ and $c(z_{1},\cdot)$ is continuous, we have, as $N\to\infty$,
$m_{N}(z_{1})=\max_{1\leq k\leq N}c(z_{1},q(k))\uparrow\max_{1\leq k<\infty}c(z_{1},q(k))=\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\quad\text{ for all }z_{1}\in\mathcal{Z}_{1}.$
Moreover, $|m_{N}(z_{1})|\leq\sup_{z_{2}\in\mathcal{Z}_{2}}|c(z_{1},z_{2})|\leq C$ for some constant $C<\infty$ independent of $z_{1}$ due to continuity of $c(\cdot,\cdot)$ and compactness of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$. Hence, by dominated convergence,
$\lim_{N\to\infty}\int m_{N}(z_{1})\,\pi_{1}^{0}(dz_{1})=\int\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi_{1}^{0}(dz_{1}).$
Therefore,
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]\geq\sup_{N\geq 1}\operatorname{\mathbb{E}}_{\pi_{N}}[c(Z_{1},Z_{2})]=\int\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi_{1}^{0}(dz_{1}).$
Combining with the upper bound (7) yields
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\pi_{1}^{0}(dz_{1})=\operatorname{\mathbb{E}}\left[\max_{z_{2}\in\mathcal{Z}_{2}}c(Z_{1},z_{2})\,\middle|\,S=0\right],$
which completes the proof.